OpenStudy (anonymous):

Does this series converge or diverge? Which method did you use? 1/(n(n+3)) w/ sigma to infinity n=1

OpenStudy (anonymous):

1/3 \[\int\limits_{1}^{infi}\] ((n+3) - n)/n(n+3)

OpenStudy (anonymous):

now you can solve the integration easily...

OpenStudy (anonymous):

if you worked this out, did you get 11/18? and do you mind also telling me if n+1/2n-1 converges? to 1/2?

OpenStudy (anonymous):

Comparison method would work here as well.

OpenStudy (anonymous):

Ratio would work too actually.

OpenStudy (anonymous):

I think.

OpenStudy (anonymous):

But comparison seems to be the easiest.

OpenStudy (anonymous):

\[ n(n+3)>n(n) \implies \frac{1}{n(n+3)} < \frac{1}{n^2} \]

OpenStudy (anonymous):

Yep that was my guess too.

OpenStudy (anonymous):

That converges obviously.

OpenStudy (anonymous):

Because the exponent is greater than 1.

OpenStudy (anonymous):

i think you can actually compute this one exactly

OpenStudy (anonymous):

It has to be an telescoping series though for that.

OpenStudy (anonymous):

it does telescope

OpenStudy (anonymous):

P-series test.

OpenStudy (anonymous):

Ohh yes it does! :P .

OpenStudy (anonymous):

Forgot about tpartial fractions.

OpenStudy (anonymous):

\[\sum\frac{1}{n(n+3)}=\frac{1}{3}\sum \frac{1}{n}-\frac{1}{n+3}\]

OpenStudy (anonymous):

The series in the first question is telescoping so they might be talking about that one? okay one more, can you help with a method to solve this one? (1/n) - (1/n+2)

OpenStudy (anonymous):

For that one I would use the integral test.

OpenStudy (anonymous):

And by the looks of it, it's going to diverge.

OpenStudy (anonymous):

that one telescopes too

OpenStudy (anonymous):

Really? It looks like it's going to diverge to me.

OpenStudy (anonymous):

if you are going to use a test for convergence, and not actually add, you have to add them up first

OpenStudy (anonymous):